OpenStudy (anonymous):

Find a formula for f o g given the functions f(x)=13x^-13 g(x)=3x^2

OpenStudy (anonymous):

I assume f(x) = 12x^2 - 13 fog = 13 [ ( 3x2)^2 - 1 ] = 13 [ 9x^4 - 1] = 117x^4 - 13

OpenStudy (anonymous):

\[f(x)=13x ^{-13}\] \[g(x)=3x ^{2}\] \[fog(x)=f(g(x))=f(3x ^{2})=13(3x ^{2})^{-13}\] \[39x ^{-26}\]